OpenStudy (anonymous):

A and B are constants such that (x+6)/(x^2 -3x) =A/x +B/(x-3). Find A-B.

Parth (parthkohli):

Again partial fractions?

Parth (parthkohli):

Start with\[\rm \large {x + 6 \over x(x - 3)} = {A \over x} + {B \over x - 3}\]Multiply both sides by \(\rm x(x - 3)\) to clear the fractions out.

OpenStudy (anonymous):

wouldn't that just make the other side more complicated?

Parth (parthkohli):

No, it'd make it easier. Do what I said.

OpenStudy (anonymous):

how would you multiply it to the other side?

OpenStudy (anonymous):

would you get x+6=Ax-3A+Bx

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

so would A=-2 B=3